Solve the quadratic equation by completing the perfect square
X squared -14x-5=0

Answer by stanbon(75887)

Solve the quadratic equation by completing the perfect square
X squared -14x-5=0
----
x^2 - 14x + ? = 5 + ?
---
x^2 - 14x + 7^2 = 5 + 7^2
---
Factor:
(x-7)^2 = 54
---
x-7 = sqrt(54) or x-7 = -sqrt(54)
---
x = 7+sqrt(54) or x = 7-sqrt(54)
====================================
Cheers,
Stan H.
===========